# Latent Reasoning in LLMs as a Vocabulary-Space Superposition

### v1

@@ Abstract
Large language models (LLMs) demonstrate strong reasoning abilities with chain-of-thought prompting, but explicit reasoning introduces substantial computational overhead. Recent work on latent reasoning reduces this cost by reasoning in latent space without explicit supervision, but performance drops significantly. Our preliminary experiments suggest that this degradation stems from the unstructured latent space, which makes fitting latent tokens difficult. To address this, we restrict the latent space to the column space of the LLM vocabulary, treating latent reasoning as a superposition over vocabulary probabilities. Once latent reasoning concludes, it collapses into an eigenstate of explicit reasoning to yield the final answer. Based on this idea, we propose Latent-SFT, a two-stage learning framework. In the first stage, we design two specialized attention masks to guide the Latent Token Encoder in generating latent tokens, allowing the LLM to produce the correct answer conditioned on them. In the second stage, the Latent Token Encoder is discarded, and the LLM is directly trained to generate these latent tokens autonomously for latent reasoning, optimized with KL and CE losses. Latent-SFT sets a new state of the art on GSM8k, matching explicit SFT performance while cutting reasoning chains by up to 4× and outperforming prior latent methods. On Math500 and AIME24, lexical probability–based latent reasoning also clearly surpasses hidden-state–based approaches. Our metrics of effective compression rate and effective global parallelism further show that latent reasoning is both the compression of a single path and the superposition of multiple paths. Our code, datasets, and models are publicly available at: https://github.com/DJC-GO-SOLO/Latent-SFT.
11footnotetext: Corresponding author.
Figure 1: Comparison of Latent and Explicit Reasoning. Latent reasoning operates in a vocabulary superposition state, completing the reasoning process and reaching the correct answer with fewer tokens. Each step of explicit reasoning can be regarded as approximating a one-hot distribution.
@@ 1 Introduction
Large language models (LLMs) have shown strong reasoning abilities (Zhao & Zhang, 2024) across diverse tasks (Cao & Zhao, 2025; Liu et al., 2025; Deng et al., 2025b; a; Wei et al., 2025a). A key driver of this success is chain-of-thought prompting (Wei et al., 2022; Xu et al., 2024; Yao et al., 2023), which solves complex problems by generating intermediate steps in natural language. However, current reasoning models (Team, 2025; DeepSeek-AI & et al., 2025) rely on long chains of tokens (Xu et al., 2025), leading to redundancy and high computational cost. Existing solutions, such as token pruning (Xia et al., 2025; Wei et al., 2025b) and reinforcement learning with length penalties (Arora & Zanette, 2025), reduce token usage but still constrain reasoning to the natural language space. In this work, we explore Latent Reasoning, a framework that enables reasoning in continuous latent space.
Unlike explicit reasoning, latent reasoning retains the last hidden state from the latent layer of the LLM, bypassing token sampling and directly feeding it back as the next-step representation. This hidden state is expected to carry richer information than a single token, effectively increasing bandwidth and enabling exploration of more reasoning paths with fewer generation steps (Zhu et al., 2025). While recent latent reasoning methods have succeeded in reducing token usage, they often suffer from substantial performance degradation (Hao et al., 2024).
To enhance latent reasoning, we analyze its challenges by comparing it with explicit reasoning. Firstly, supervision is difficult. Explicit reasoning leverages natural language as an anchor, enabling supervision through annotation or sampling data, with labels transferable across models. Latent reasoning, however, lacks natural labels; even if latent tokens are generated, they are often model-specific and hard to reuse. Secondly, learning is difficult. Explicit labels are discrete distributions, easy to optimize, and aligned with pre-training objectives. By contrast, latent tokens are typically the hidden states of the LLM’s last layer—high-dimensional, unconstrained, and usually optimized via MSE or cosine distance, which diverges from the pre-training task. In summary, effective latent reasoning requires a principled definition of latent tokens that supports model learning.
In Section 3, we present preliminary experiments leading to two key observations. Observation 1: the hidden state distribution of the LLM’s last layer is entirely inconsistent with the token embedding distribution. Since LLMs are trained with token embeddings as inputs, directly feeding hidden states introduces out-of-distribution values, preventing the model from interpreting their semantics and degrading performance. Observation 2: the effective rank of token embeddings in current LLMs is not full, indicating that the semantic space is inherently low-dimensional. Without constraining latent tokens, the model risks losing semantic consistency and struggles to learn meaningful representations. Based on these findings, we define latent tokens as elements within the column space of token embeddings. This ensures that latent and explicit tokens share consistent distributions and reside in the same low-rank space, thereby addressing the two central challenges revealed by our experiments.
Building on our definition of latent tokens, we introduce Latent-SFT, a two-stage learning framework. Stage 1: Latent token generation. We construct an encoder–decoder architecture using the same LLM, where the encoder is forced to produce latent tokens that enable the decoder to generate the correct answer. Specifically, the Latent Token Induction Mask (LTIM) constrains special tokens in the encoder to attend only to a sub-segment of the explicit reasoning chain. Latent tokens are then derived by linearly combining the vocabulary probabilities associated with these tokens. The Latent Token Supervision Mask (LTSuM) further enforces that each latent token decodes the subsequent explicit reasoning segment and the final answer. Together, these mechanisms ensure that latent tokens remain semantically compact, compatible, and correct. Stage 2: Latent token learning. The decoder learns to generate latent tokens independently. Specifically, KL loss is used to learn latent tokens, while CE loss is applied to optimize the final explicit answer.
As illustrated in Figure 1, our model performs latent reasoning in the column space of the vocabulary, analogous to wavefunction superposition in Hilbert space. Once reasoning is complete, entropy collapses into an eigenstate, producing the final natural language answer. Experiments confirm that Latent-SFT surpasses other latent baselines and achieves performance comparable to explicit SFT models on low-difficulty mathematical reasoning datasets. On high-difficulty datasets, latent tokens defined via soft embeddings significantly outperform those defined by hidden states. Further analyses reveal that its latent reasoning process is not merely a compression of a single reasoning chain but also a superposition of multiple chains.
@@ 2 Related work
Our work centers on latent inference, and we categorize related research into two groups: training-based methods and training-free methods.
Training-based Latent Reasoning.
COCONUT (Hao et al., 2024) first proposed using the last hidden state as the input for the next token without applying the language model head, laying the foundation for latent reasoning. However, it relied on a fixed number of latent steps and did not address the misalignment between hidden (latent) states and token embeddings, leading to weak performance. CCOT (Cheng & Durme, 2024) introduced continuous contemplation tokens and extended this idea to variable-length latent reasoning, achieving modest gains. CODI (Shen et al., 2025) improved alignment by adding an MLP to map hidden states into the token embedding space, and further adopted a teacher–student distillation strategy, forcing the latent model to match hidden states across all layers up to the final token of the explicit model—significantly boosting performance. PCCOT (Wu et al., 2025b) enhanced CODI by applying Jacobian iteration to improve both training and inference efficiency. CoLaR (Tan et al., 2025) advanced this line by introducing a latent head to predict compressed embedding distributions, maintaining alignment between latent and token spaces. Combined with reinforcement learning, CoLaR achieved state-of-the-art latent reasoning, though still trailing explicit SFT. Unlike these approaches, we restrict latent reasoning to the column space of the LLM vocabulary, directly addressing misalignment between latent and token spaces.
Training-free Latent Reasoning.
Several works have explored training-free latent reasoning, including Soft Thinking (Zhang et al., 2025) and Mixture-of-Input (MoI) (Zhuang et al., 2025). Like our approach, they use next-token probability distributions to linearly combine token embeddings, thereby linking hidden and embedding spaces. However, Wu et al. (2025a) showed that this remains a form of greedy decoding, which limits performance. In contrast, our method can be seen as a trained variant of Soft Thinking, enabling the model to explore multiple reasoning paths.
@@
3 Preliminary Experiments: Definition of Latent Token
To gain a deeper understanding ofing of the latent reasoning process, we conducted two preliminary experiments addressing the core question of what the representation of latent tokens should be.
3.1 Distributional Discrepancy Between Final-Layer Hidden States and Token Embeddings
Most existing studies (Shen et al., 2025; Wu et al., 2025b) follow the COCONUT design, using the last-layer hidden state directly as the representation of the next latent token rather than sampling from the explicit token embedding matrix. However, a statistical analysis of the last-layer hidden states and token embeddings in the representative LLMs reveals the following observation.
Observation 1. The distribution of last-layer hidden states in the LLM is entirely inconsistent with that of the token embeddings. As shown in Figure 2 and Figure 7, the distributions of hidden states and token embeddings after PCA dimensionality reduction differ substantially. Statistically, their means, variances, and inter-distribution distances are all large. From a modeling perspective, shallow parameters (e.g., RMSNorm) encountered during pre-training are exposed only to the token embedding distribution. Feeding vectors from an unfamiliar distribution into the model can thus induce distribution shift or other unpredictable effects, which is one of the reasons why methods such as COCONUT perform poorly.
3.2 Non-Full Effective Rank of Token Embeddings in Large Language Models
@@
Building on Observation 1, we posit that the latent token representations should ideally align with the distributional geometry of the token embeddings. To examine this property, we analyze the singular value spectra of token embedding matrices across different LLMs.
@@
Observation 2. The Effective Rank of LLM Token Embeddings Is Not Full. The resulting curves in Figure 3 exhibit a rapid decay rather than a flat plateau, indicating that only a subset of singular directions contributes substantially to the embedding space. The effective rank of word embeddings is far below the theoretical maximum, indicating that although the embedding matrix has high nominal dimensionality, it occupies only a low-dimensional subspace. Thus, the semantic space of LLMs can be regarded as inherently low-dimensional.
@@
These findings suggest that latent tokens should not be treated as unconstrained hidden vectors, but rather as structured elements residing within the embedding subspace. Guided by this insight, we formally define latent tokens.
Definition 1 (Soft Embedding). Formally, a latent token $z\in\mathbb{R}^{d}$ can be expressed as a linear combination of token embeddings:
$z=\sum_{i=1}^{V}\alpha_{i}e_{i},\quad\mathrm{with}\ \alpha_{i}\in\mathbb{R},\\ e_{i}\in\mathbb{R}^{d},$ (1)
@@
where $\{e_{i}\}_{i=1}^{V}$ denotes the embedding vectors of the vocabulary, and the coefficients $\{\alpha_{i}\}$ determine the semantic mixture. This formulation ensures that latent tokens remain aligned with the intrinsic semantic geometry of the LLM’s vocabulary space, providing a principled interpretation of latent tokens as interpolation within the embedding manifold.
4 Latent-SFT: Equipping Large language Models with Latent Reasoning
With Definition 1, we introduce Latent-SFT to bring latent reasoning into LLMs. As a prerequisite, latent tokens must be generated (Definition 1), since they provide the labels needed for supervision. Thereafter, the model is trained, in analogy to explicit reasoning, to generate these latent tokens and perform latent reasoning. The overall training framework is illustrated in Figure 4.
@@ 4 Latent-SFT: Equipping Large language Models with Latent Reasoning
Figure 4: Overview of Latent-SFT. The training process consists of two phases: (a) Generating Latent Tokens via Induction–Supervision Masking. In the mask sub-figure, components such as $Q$, $S_{i}$ and Answer (which may span multiple tokens) are implemented via a standard autoregressive attention mechanism. The Latent Token Encoder shares both structure and initialization with the base LLM. (b) Training the LLM to Autonomously Generate Latent Tokens. In this phase, the LLM is trained to perform latent reasoning independently, using a weighted combination of KL loss and CE loss.
4.1 Generating Latent Tokens via Induction-Supervision Masking
Latent tokens are designed to be: (1) semantic compactness, replacing multiple explicit tokens; (2) semantic compatibility, constrained to the semantic space of explicit tokens (Section 3); and (3) semantic correctness, capable of producing the right answer. We implement these properties within an encoder–decoder framework through three key steps: information compression, representation alignment, and supervised decoding.
@@ 4.1 Generating Latent Tokens via Induction-Supervision Masking
We insert special tokens into the explicit reasoning sequence to segment and capture intermediate reasoning information. Given an input $X=[Q,\text{\textless think\textgreater},T,\text{\textless/think\textgreater}$], we define a segmentation function $\mathrm{Seg}(\cdot)$ that divides $T$ (explicit reasoning sequence) into $N$ subsegments, i.e., $T=\{S_{i}\}_{i=1}^{N}$. This segmentation can be performed either by a fixed token length (compression ratio) $r$ or by grouping semantically continuous thoughts. After each $S_{i}$ , we insert a special token $L_{i}$ and apply the Latent Token Induction Mask (LTIM) to restrict each $L_{i}$ to attend only to the explicit reasoning information preceding it (see Figure 4). This design ensures that the semantic content encoded by latent tokens grows with reasoning completeness, facilitating subsequent model learning.
@@
Given the inserted special tokens $\{L_{i}\}_{i=1}^{N}$, we first obtain their hidden states $h_{i}\in\mathbb{R^{d}}$ from the (LTIM-masked) model. To align each latent token with the LLM’s vocabulary manifold, we project $h_{i}$ onto the vocabulary simplex and then reconstruct its embedding as a mixture of token embeddings, in accordance with Definition 1. Concretely, let $E=[e_{1},...,e_{V}]\in\mathbb{R}^{d\times V}$ denote the (input) embedding matrix and $W\in\mathbb{R}^{d\times V}$ the LM head (often $W=E$) under weight tying). We compute vocabulary logits and a probability vector
$\bm{\ell}_{i}=W^{\!\top}\mathbf{h}_{i},\qquad\bm{\alpha}_{i}=\mathrm{softmax}\!\left(\bm{\ell}_{i}/T\right)\in\Delta^{V-1},$ (2)
where $T>0$ is an optional temperature. The representation of the $i$-th latent token is then obtained by linear interpolation within the embedding space:
$\mathbf{z}_{i}\;=\;E\,\bm{\alpha}_{i}\;=\;\sum_{v=1}^{V}\alpha_{i,v}\,\mathbf{e}_{v}.$ (3)
which instantiates Definition 1 (Soft Embedding) with coefficients given by the model-implied vocabulary distribution. This construction guarantees $z_{i}$ lies in the span (and, with softmax, the convex hull) of $\{e_{v}\}_{v=1}^{V}$, thereby enforcing semantic compatibility: latent tokens remain confined to the intrinsic geometry of the LLM’s vocabulary space. In training Stage 2, $\alpha_{i}$ serves as the supervision signal for latent positions (KL term), while explicit positions are optimized with CE; optionally, one may prune $\alpha_{i}$ to top-$k$ and renormalize to reduce computation without departing from the embedding manifold.
Supervised Decoding Process (Semantic Correctness).
Given the latent representations $\{z_{i}\}_{i=1}^{N}$ from the alignment step, we use each $z_{i}$ to decode the content from the (i+1i\!+\!1)-th explicit reasoning subsegment through to the final answer. Let the prefix up to step $i$ be $\Pi_{i}\!=\![Q,\text{\textless think\textgreater},z_{1},S_{2},z_{2},...,z_{i}]$. Under the Latent Token Supervision Mask (LTSuM), tokens in $\mathcal{Y}_{i}\!=\![S_{i+1},...,S_{N},\text{\textless/think\textgreater},\mathrm{Answer}]$ are decoded conditioned only on $\Pi_{i}$ (in particular on $Q,\text{\textless think\textgreater},z_{1},...,z_{i}]$), while attention to future latent tokens and to past explicit reasoning steps is blocked. Denoting by $\mathcal{J}_{i}$ the index set of tokens in $\mathcal{Y}_{i}$, the supervised decoding objective is

### v2

# LLM Latent Reasoning as Chain of Superposition

@@ Abstract
Latent reasoning offers a computation-efficient alternative to Chain-of-Thought but often suffers from performance degradation due to distributional misalignment and ambiguous chain definitions. Ideally, latent reasoning should function as a superposition of multiple reasoning paths. To realize this, we introduce Latent-SFT, a unified framework addressing challenges at three levels: token, chain, and learning. First, we define the Latent-Vocab to constrain hidden states within the pre-trained vocab-space. Second, we construct the Latent-Chain via Induction-Supervision Masking to ensure semantic compactness and sufficiency. Third, we employ Latent-Optim with stochastic Gumbel-Softmax to guide the model toward generalizable solutions. Empirical results demonstrate that Latent-SFT consistently outperforms explicit SFT across six mathematical benchmarks (e.g., GSM8k, AIME24) while achieving a 2.7x to 5.5x reduction in reasoning length. Analysis confirms that our method effectively captures a superposition of diverse reasoning trajectories rather than merely compressing a single path. The models and datasets of our method are available in a repository: https://anonymous.4open.science/r/Latent-SFT-CF17.
Machine Learning, ICML
Figure 1: Comparison of Latent and Explicit Reasoning. Latent reasoning operates as a superposition state, where each step carries a higher entropy. Therefore, it can complete the reasoning process and derive the correct answer using fewer tokens.
@@ 1 Introduction
Large language models (LLMs) have shown strong reasoning abilities (Zhao and Zhang, 2024) across diverse tasks (Cao and Zhao, 2025; Liu et al., 2025; Deng et al., 2025b, a; Wei et al., 2025). A key driver of this success is chain-of-thought prompting (Wei et al., 2022; Xu et al., 2024; Yao et al., 2023), which solves complex problems by generating intermediate steps in natural language. However, current reasoning models (Team, 2025; DeepSeek-AI, 2025) rely on long chains of tokens (Xu et al., 2025), leading to redundancy and high computational cost. Latent Reasoning is a promising paradigm that significantly reduces the length of reasoning chains by shifting computation to a continuous latent space. This paradigm theoretically enhances information bandwidth and exploration efficiency (Zhu et al., 2025), yet current methods frequently suffer from severe performance degradation compared to explicit reasoning (Hao et al., 2024).
Ideally, latent reasoning should function as a superposition of multiple reasoning paths, enabling the model to explore diverse logical trajectories simultaneously within a latent space. However, realizing this potential is impeded by three hierarchical flaws in current methodologies. First, at the latent token level, existing approaches rely on raw, unconstrained hidden states (Hao et al., 2024; Cheng and Durme, 2024). These high-dimensional vectors suffer from severe distributional shifts relative to the model’s vocab-space, making them difficult for the model to interpret. Second, at the latent chain level, current definitions are ambiguous; they typically assume the latent chain need only decode the final answer (Shen et al., 2025). This ”black box” assumption neglects intermediate semantic fidelity, failing to capture the step-by-step logic required for complex deduction. Third, at the latent learning level, optimization is inefficient. Models are updated solely via indirect backpropagation from the final explicit supervision (Hao et al., 2024), lacking the step-by-step guidance necessary to learn stable reasoning trajectories.
To systematically address these challenges, we introduce Latent-SFT, a comprehensive framework that unifies token-level representation, chain-level definition, and learning strategies. For token-level representation, we formally define the Latent-Vocab as a superposition within the vocab-space (Section 3). This formulation aligns latent tokens with the model’s inherent vocabulary distribution. For the chain-level definition, we introduce the Latent-Chain (Section 4), constructed via Induction-Supervision Masking to satisfy three critical semantic properties. First, the Latent Token Induction Mask (LTIM) enforces semantic compactness by compressing explicit reasoning segments into dense representations. Second, these representations are formalized as a Latent-Vocab—a superposition within the vocab-space—ensuring semantic compatibility with the model’s semantic manifold. Third, the Latent Token Supervision Mask (LTSuM) guarantees informational sufficiency by mandating that these tokens can reconstruct subsequent reasoning steps. Finally, regarding the learning strategy, to enable the LLM to autonomously generate these latent chains, we propose Latent-Optim. Within this framework, we address the static nature of the constructed latent chains by employing stochastic Gumbel-Softmax optimization. This technique injects Gumbel noise to approximate sampling, acting as an implicit Hessian regularizer (see Theorem 1 in Appendix C) that guides the model toward flatter, more generalizable minima.
As illustrated in Figure 1, our model realizes the ideal latent reasoning via a chain of superposition analogous to a quantum wavefunction. Upon reasoning convergence, the system undergoes entropy collapse, precipitating the latent superposition into a discrete linguistic eigenstate. Empirical results demonstrate that Latent-SFT consistently outperforms explicit SFT models across six diverse mathematical reasoning benchmarks. Notably, it achieves a 2.7x to 5.5x reduction in reasoning chain length while surpassing prior latent reasoning approaches. Further analysis reveals that its latent reasoning is not merely a compression of a single chain, but rather a superposition of multiple reasoning chains.
@@ 2 Related Work
Current research remains entangled in debates over the reasonable representation of latent tokens, while the critical dimensions of latent chain definition and optimization strategies have been largely overlooked. Consequently, the definition of the latent token serves as the primary differentiator in this field. We categorize existing research based on this pivotal factor.
@@
Unconstrained Representations.
Pioneering approaches, such as COCONUT (Hao et al., 2024) and CCOT (Cheng and Durme, 2024), adopted a direct approach by utilizing the last-layer hidden states as latent tokens. While computationally straightforward, these methods rely on raw, high-dimensional vectors that are unconstrained by the model’s vocabulary distribution. Although subsequent works like CODI (Shen et al., 2025) and PCCOT (Wu et al., 2025b) attempted to mitigate this misalignment via MLP mapping or distillation, they fundamentally remain bound to the continuous hidden state paradigm. Consequently, without strict constraints aligned with the pre-trained semantic manifold, these methods often face optimization instability and limited performance gains compared to explicit reasoning.
@@
Constrained Representations.
To address the distributional gap, recent studies have introduced structural constraints. CoLaR (Tan et al., 2025) represents a step forward by modeling latent tokens via Gaussian distributions, achieving improved alignment through a dedicated latent head. Similarly, methods like Soft Thinking (Zhang et al., 2025) and Mixture-of-Input (MoI) (Zhuang et al., 2025) attempt to bridge the gap by constructing latent tokens as linear combinations of vocabulary embeddings. However, the performance gains of these embedding-merge approaches remain marginal, and crucially, they fail to reduce the reasoning chain length. As analyzed by (Wu et al., 2025a), these training-free methods operate essentially as greedy decoding in a latent space. Consequently, they are incapable of realizing the ideal state of latent reasoning—namely, the learned superposition of multiple reasoning trajectories. In contrast, we formally define the Latent-Vocab and Latent-Chain to enable simultaneous multi-path exploration, thereby significantly shortening reasoning chains while surpassing existing baselines. Given that training-free methods neither involve parameter optimization nor achieve the reasoning length reduction central to our contribution, we exclude them from our primary baselines to focus on training-based methods that deliver compression of the reasoning chain.
3 Latent-Vocab: Superposition of Vocab-Space
Unlike prior studies that rely on raw, unconstrained hidden states, we conducted two preliminary experiments addressing the fundamental question: what is the reasonable representation for latent tokens?
3.1 The Distributional Discrepancy in Vector Spaces
Most existing studies (Shen et al., 2025; Wu et al., 2025b) follow the COCONUT design, using the last-layer hidden state directly as the representation of the next latent token rather than sampling from the explicit token embeddings. However, a statistical analysis of these vector spaces in representative LLMs reveals a critical discrepancy.
Observation 1. Significant distributional divergence exists between the last-layer hidden states and the token embedding space. As illustrated in Figure 2 and Figure 8, PCA projections demonstrate a substantial separation between the manifolds of hidden states and token embeddings. Quantitative analysis confirms this divergence, showing significant disparities in first and second-order statistics (mean and variance), as well as large inter-distribution distances. From a model optimization perspective, shallow parameters (e.g., RMSNorm) are calibrated exclusively on the token embedding distribution during pre-training. Consequently, introducing vectors from an alien distribution (the raw hidden states) can precipitate distribution shift and training instability, offering a potential explanation for the performance limitations observed in methods such as COCONUT.
3.2 The Low Effective Rank of Vocab-Space
@@
Building on Observation 1, we posit that latent token should ideally preserve the distributional geometry of vocab-space. To investigate the intrinsic dimensionality of this space, we analyze the singular value spectra of token embedding matrices across various LLMs.
@@
Observation 2. The Token Embedding Matrix exhibits a non-full effective rank. The singular value decay curves presented in Figure 3 exhibit a precipitous drop rather than a gradual plateau, indicating that the vocab-space is dominated by a small subset of singular directions. Despite the high nominal dimensionality of the embedding matrix, its effective rank is considerably lower than the theoretical maximum. This suggests that the semantic manifold of LLMs is inherently low-dimensional.
@@
These findings imply that latent tokens should not be modeled as unconstrained continuous vectors, but rather as structured elements constrained to the embedding subspace. Motivated by this geometric insight, we propose the following formal definition.
Definition 1 (Superposition Embedding). We define a latent token $z\in\mathbb{R}^{d}$ as a vector residing within the linear span of the vocab-space. Formally, let $\mathbf{E}=[e_{1},e_{2},\dots,e_{V}]\in\mathbb{R}^{d\times V}$ denote the token embedding matrix, where $e_{i}\in\mathbb{R}^{d}$ represents the embedding vector for the $i$-th token in a vocabulary of size $V$. A latent token $z$ is expressed as a linear combination of these embeddings:
$z=\mathbf{E}\bm{\alpha}=\sum_{i=1}^{V}\alpha_{i}e_{i},$ (1)
@@
where $\bm{\alpha}=[\alpha_{1},\dots,\alpha_{V}]^{\top}\in\mathbb{R}^{V}$ represents the coefficient vector governing the semantic mixture. This formulation ensures that latent tokens remain aligned with the intrinsic semantic geometry of the LLM’s vocab-space.
@@
Figure 4: Overview of Latent-SFT. The entire training process consists of two phases: (a) Generating Latent Chains via Induction–Supervision Masking. In the mask sub-figure, components such as $Q$, $S_{i}$ and Answer (which may span multiple tokens) are implemented via a standard autoregressive attention mechanism. The Latent Token Encoder shares both structure and initialization with the base LLM. (b) Training the LLM to Autonomously Generate Latent Chains. In this phase, the LLM is trained to perform latent reasoning independently, using a weighted combination of KL loss and CE loss.
4 Latent-Chain: Forming Latent Chain via Induction-Supervision Masking
Having established the definition of the Latent-Vocab, we address the next core question: what are the requisite properties of the higher-level Latent-Chain? Unlike prior approaches that treat the latent chains as an ambiguous “black box” optimized solely for final answer reconstruction, thereby neglecting intermediate semantic fidelity, we aim to capture the step-by-step logic essential for latent reasoning by postulating that the Latent-Chain must satisfy three key semantic criteria: (1) Semantic Compactness, compressing multiple explicit tokens into a single latent representation; (2) Semantic Compatibility, adhering to the intrinsic semantic space of explicit tokens (as per Section 3); and (3) Semantic Sufficiency, retaining sufficient information to deduce the correct answer. As shown in the Figure 4, we implement these properties within a encoder-decoder framework through three key steps: information compression, representation alignment, and supervised decoding.
@@ 4 Latent-Chain: Forming Latent Chain via Induction-Supervision Masking
We insert special tokens into the explicit reasoning sequence to segment and capture intermediate reasoning information. Given an input $X=[Q,\texttt{<think>},C,\texttt{</think>}]$, we define a segmentation function $\mathrm{Seg}(\cdot)$ that partitions the explicit reasoning chain $C$ into $N$ subsegments, denoted as $C=\{S_{i}\}_{i=1}^{N}$. This segmentation can be governed either by a fixed compression ratio $r$ or by semantic boundaries. Following each subsegment $S_{i}$, we insert a special latent placeholder $L_{i}$. Crucially, we apply a Latent Token Induction Mask (LTIM) to restrict the attention of $L_{i}$ exclusively to the preceding explicit reasoning context (see Figure 4). This design ensures that the semantic content encoded by each latent token represents a cumulative summary of the reasoning process up to that point.
@@
For each inserted placeholder $L_{i}$, we extract its hidden state $\mathbf{h}_{i}\in\mathbb{R}^{d}$ from the model under LTIM masking. To strictly enforce Semantic Compatibility, we project $\mathbf{h}_{i}$ onto the vocabulary simplex and reconstruct it as a mixture of token embeddings. Let $\mathbf{E}=[\mathbf{e}_{1},\dots,\mathbf{e}_{V}]\in\mathbb{R}^{d\times V}$ denote the token embedding matrix and $\mathbf{W}\in\mathbb{R}^{d\times V}$ be the language modeling head (where $\mathbf{W}=\mathbf{E}$ under weight tying). We compute the vocabulary logits and the resulting probability distribution as:
$\bm{\ell}_{i}=\mathbf{W}^{\!\top}\mathbf{h}_{i},\quad\bm{\alpha}_{i}=\mathrm{softmax}\!\left(\frac{\bm{\ell}_{i}}{\tau}\right)\in\Delta^{V-1},$ (2)
where $\tau>0$ is the temperature parameter. The final latent representation is then obtained via linear interpolation:
$\mathbf{z}_{i}=\mathbf{E}\bm{\alpha}_{i}=\sum{v=1}^{V}\alpha_{i,v}\mathbf{e}_{v}.$ (3)
This operation instantiates Definition 1, guaranteeing that $\mathbf{z}_{i}$ resides within the convex hull of the embedding space $\{\mathbf{e}_{v}\}_{v=1}^{V}$. Consequently, latent tokens remain confined to the intrinsic geometry of the LLM’s vocabulary manifold. During the subsequent optimization phase, the distribution $\bm{\alpha}_{i}$ serves as the supervision target (via KL divergence) for the latent positions.
Supervised Decoding Process (Semantic Sufficiency).
Given the latent representations $\{\mathbf{z}_{i}\}_{i=1}^{N}$ from the alignment step, we use each $\mathbf{z}_{i}$ to decode the content from the (i+1i\!+\!1)-th explicit reasoning subsegment through to the final answer. Let the prefix up to step $i$ be $\Pi_{i}=[Q,\texttt{<think>},\mathbf{z}_{1},S_{2},\mathbf{z}_{2},\dots,\mathbf{z}_{i}]$. Under the Latent Token Supervision Mask (LTSuM), tokens in $\mathcal{Y}_{i}=[S_{i+1},\dots,S_{N},\texttt{</think>},\mathrm{Answer}]$ are decoded conditioned only on $\Pi_{i}$ (in particular on $Q,\texttt{<think>},\mathbf{z}_{1},...,\mathbf{z}_{i}]$), while attention to future latent tokens and to past explicit reasoning steps is blocked. Denoting by $\mathcal{J}_{i}$ the index set of tokens in $\mathcal{Y}_{i}$, the supervised decoding objective is